# Photometric stellar masses for galaxies in DESI Legacy Imaging Surveys

### v1

@@ 2.4 GALFIT modeling
Table 1: Parameters, RMS values, and the number of discarded outlying galaxies ($N_{\text{out}}$) of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes in different bands.
@@ 3.1 The sample
Integrated magnitudes were extracted from the Legacy Surveys DR10 by the procedure in the previous section, Sect. 2, and corrected for the Galactic extinction as described in Sect. 4. Absolute magnitudes, $M_{g}$, $M_{r}$, and $M_{z}$, were computed using the same mean redshift-independent distances that were used in S4G to derive the stellar masses. The dataset with all related parameters is available at https://doi.org/10.5281/zenodo.14253992.
@@ 3.4 Including other quantities
Table 2: Parameters, RMS values, and the number of discarded outlying galaxies ($N_{\text{out}}$) of different fitting formulas using total magnitudes in $g$ and $r$ bands. The reference S4G stellar masses are computed according either Meidt et al. (2014) or Querejeta et al. (2015) prescriptions.
@@ 3.5 Other S4G mass calibrations
Table 3: Parameters and RMS values of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes computed from $M_{26}$ and $M_{\text{TOT}}$ magnitudes in SGA-2020.
@@ 3.8 Rest-frame colors
As the $g$ and $r$ filters of the Legacy Surveys instruments were designed to be similar to SDSS filters, we employed, for these filters, the $K$-corrections calculator [^5] http://kcor.sai.msu.ru/ (Chilingarian et al., 2010; Chilingarian & Zolotukhin, 2012), which was designed for SDSS colors. Redshifts of galaxies were calculated from heliocentric radial velocities from radio measurement provided by S4G catalog. For the forty galaxies in our sample that do not have the radial velocities listed in S4G, we supplied the value from the HyperLeda database [^6] http://leda.univ-lyon1.fr/ (Makarov et al., 2014).
@@ 4 Script for stellar mass estimates
We provide a Python script named photomass_ls.py, which is publicly available at: https://github.com/PidiGalaxies/photomass.
@@
The photomass_ls.py script provides additional optional functionalities, including photometry in the $z$ and $i$ bands of Legacy Surveys images and $K$ correction for rest-frame colors of galaxies with redshifts up to 0.5. An example of script usage is provided in Appx. C, with additional examples available in the user manual at https://github.com/PidiGalaxies/photomass. The full manual, with additional examples, is available at https://github.com/PidiGalaxies/photomass.
@@ Appendix A Additional tested formulas
Table 4: Parameters and RMS values of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes, Sérsic indexes, and axis ratios in different bands.

### v3

@@ 2.4 GALFIT modeling
Table 1: Fitted formulas  [^4] Parameters, RMS values, and the number of discarded outlying galaxies ($N_{\text{out}}$) of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes in different bands.
@@ 3.1 The sample
Integrated magnitudes were extracted from the Legacy Surveys DR10 by the procedure in the previous section, Sect. 2, and corrected for the Galactic extinction as described in Sect. 4. Absolute magnitudes, $M_{g}$, $M_{r}$, and $M_{z}$, were computed using the same mean redshift-independent distances that were used in S4G to derive the stellar masses. The dataset with all related parameters is available at Zenodo [^5] https://doi.org/10.5281/zenodo.14253992 and CDS.
@@ 3.4 Including other quantities
Table 2: Fitted formulas with alternative calibrations  [^7] Parameters, RMS values, and the number of discarded outlying galaxies ($N_{\text{out}}$) of different fitting formulas using total magnitudes in $g$ and $r$ bands. The reference S4G stellar masses are computed according either Meidt et al. (2014) or Querejeta et al. (2015) prescriptions.
@@ 3.5 Other S4G mass calibrations
Table 3: Fitted formulas for SGA-2020 magnitudes  [^8] Parameters and RMS values of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes computed from $M_{26}$ and $M_{\text{TOT}}$ magnitudes in SGA-2020.
@@ 3.8 Rest-frame colors
As the $g$ and $r$ filters of the Legacy Surveys instruments were designed to be similar to SDSS filters, we employed, for these filters, the $K$-corrections calculator [^9] http://kcor.sai.msu.ru/ (Chilingarian et al., 2010; Chilingarian & Zolotukhin, 2012), which was designed for SDSS colors. Redshifts of galaxies were calculated from heliocentric radial velocities from radio measurement provided by S4G catalog. For the forty galaxies in our sample that do not have the radial velocities listed in S4G, we supplied the value from the HyperLeda database [^10] http://atlas.obs-hp.fr/hyperleda/ (Makarov et al., 2014).
@@ 4 Script for stellar mass estimates
We provide a Python script named photomass_ls.py, which is publicly available at GitHub [^12] https://github.com/PidiGalaxies/photomass.
@@
The photomass_ls.py script provides additional optional functionalities, including photometry in the $z$ and $i$ bands of Legacy Surveys images and $K$ correction for rest-frame colors of galaxies with redshifts up to 0.5. An example of script usage is provided in Appx. C.l The full manual, with additional examples, is available at GitHub [^14] https://github.com/PidiGalaxies/photomass.
@@ 6 Conclusions
7 Data availability
The stellar mass estimates, together with all basic and fitted parameters, are available in electronic form at Zenodo [^16] https://doi.org/10.5281/zenodo.14253992, and and at the CDS via anonymous FTP at cdsarc.u-strasbg.fr (130.79.128.5) or via the web interface [^17] http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Appendix A Additional tested formulas
Table 4: Fitted formulas with additional parameters  [^18] Parameters and RMS values of different fitting formulas for calculating log($M_{*}$[M]) using galaxy total magnitudes, Sérsic indexes, and axis ratios in different bands.